OpenStudy (anonymous):

I WILL GIVE A MEDAL ! A store has a $489.99 item on sale for 50% off, plus an additional 20% off. What is the sale price you would pay on this item? A: $195.99 B: $145.00 C: $145.00 D: $194.99 Yes the teacher did put the same number twice, i don't know why.

OpenStudy (anonymous):

@nphuongsun93

OpenStudy (anonymous):

Do the same like the previous one.

OpenStudy (lexilove13):

A. 195.99 is your answer. Because 489.99*0.5 = 244.995* 0.8 = 195.996